# Finite-cutoff Holographic Thermodynamics

### v1

@@ Abstract
The thermodynamics of AdS black holes has been extensively studied following the discovery of the Hawking-Page phase transition in Schwarzschild-AdS black holes Hawking and Page (1983b). Phase structures analogous to the van der Waals liquid–gas system have been identified for charged Reissner-Nordström AdS black holes, exhibiting first-order phase transitions and critical phenomena Chamblin et al. (1999a, b); Kubiznak and Mann (2012). Recently, significant advances have been achieved in AdS black hole thermodynamics by treating the cosmological constant as a variable, resulting in phenomena such as fluid/superfluid phase transitions Hennigar et al. (2017), black hole microstructures Wei et al. (2019), and more Caldarelli et al. (2000); Kastor et al. (2009); Dolan (2011a, b, c); Cvetic et al. (2011); Lu et al. (2012); Xiao et al. (2024); Frassino et al. (2023). The study of thermodynamics within this extended phase space has developed into a subdiscipline known as black hole chemistry Kubiznak et al. (2017).
@@
on the conformal boundary at spatial infinity Ahmed et al. (2023a); Gong et al. (2023); Ahmed et al. (2023b). Here, $M,T,S,V,P$ respectively represent the bulk mass, temperature, entropy, thermodynamic volume, and thermodynamic pressure, while $\bar{E},\bar{T},\bar{S},\bar{V},\bar{P}$ denote their boundary counterparts. Note that there is an extra pair $(c,\mu)$ of conjugate variables, with $c$ being the central charge and $\mu$ its conjugate chemical potential, relative to the bulk. As pointed out in Visser (2022), the internal energy of the large-$N$ $SU(N)$ gauge theories with conformal symmetry depends not only on the extensive quantities $\bar{S}$ and $\bar{V}$, but also on the intensive quantity $c\sim N^{2}$. This is consistent with the scenario of working in the extended phase space, where
$P=\frac{3}{8\pi\mathrm{G}\ell^{2}}$ (7)
relates the AdS radius to the thermodynamic pressure, with Newton’s constant $\mathrm{G}$. Variation of the pressure is equivalent to varying central charge and volume of the CFT. The standard extrapolated holographic dictionary reads Witten (1998b); Gubser et al. (1996); Verlinde (2000); Savonije and Verlinde (2001)
@@
Finite-cutoff holographic thermodynamics: bulk side—The $T^{2}$ deformation of the three-dimensional CFT with large central charge is holographically dual to the SAdS black hole with a finite Dirichlet radial cutoff. In the framework of finite-cutoff holography, there is an agreement between the dressed energy spectrum (15) of the deformed field theory and the quasilocal energy of the SAdS black hole evaluated within the region $r<r_{c}$, given by Brown et al. (1994); Hartman et al. (2019)
@@
Phase transitions—In the cutoff SAdS bulk, the free energy is defined as $G=\mathcal{E}-TS$; setting $M=0$ in (21) yields that of the cutoff AdS vacuum. As in the standard SAdS scenario, a Hawking–Page phase transition occurs: as the temperature increases, the system transits from the cutoff vacuum AdS to the cutoff SAdS black hole. On the deformed CFT side, comparing the free energy given by $\bar{G}=\bar{\mathcal{E}}-\bar{T}\bar{S}$ with that of the deformed vacuum reveals a confinement/deconfinement phase transition. Figure 1 shows the coexistence curve traced out by the phase transition points. Due to holographic duality, the $\bar{T}$ vs. $\bar{S}$ and $T/\omega$ vs. $S$ curves coincide and are plotted together. Varying $A_{c}$ or $\lambda$ amounts to changing the UV cutoff scale of the deformed CFT and selecting a family of $T^{2}$-deformed CFTs. This redefines the corresponding holographic RG flow starting from $A_{c}$ and also redefines the RG flow of the deformed CFT starting from a new UV scale. Note the deformation flow here is not the traditional RG flow; rather it defines a family of effective field theories without coarse-graining. The point $X_{0}$ marks the confinement-deconfinement transition point in the undeformed CFT or the Hawking–Page phase transition point in SAdS without a cutoff. As the deformation parameter $\lambda$ increases (reducing the cut-off $A_{c}$), the curve flows from $X_{0}\to X_{1}\to X_{2}\to X_{3}$, forming a teardrop-shaped Rupert coexistence curve. The region above the segment $X_{0}X_{1}X_{2}$ corresponds to the deconfinement (bulk black hole) phase, while the region below corresponds to the confinement (AdS vacuum) phase. As $\lambda$ increases further, the curve extends as entropy decreases back to $X_{0}$. The segment $X_{2}X_{0}$ separates the deconfinement (black hole) phase above from the confinement (AdS vacuum) phase below. The loop $X_{0}X_{1}X_{2}X_{0}$ encloses an island where the phase varies under the deformation flow, surrounded by an exterior region of fixed phase. Within the Rupert curve, up to three deformed CFTs or cut-off SAdS configurations may share the same phase transition temperature, with one deformed CFT (or cut-off SAdS) matching the seed CFT’s (or original SAdS’s) phase transition temperature and entropy.
Figure 1: Coexistence curve of the boundary confinement/deconfinement and bulk Hawking-Page phase transitions with with $c=4,\bar{V}=20$.
@@
Based on our results, several directions for future research can be pursued. Our analysis can be readily extended to holographic thermodynamics for a CFT in general dimensions, including $d=2$, where $T\bar{T}$ holographic thermodynamics can be constructed. The framework can also be applied to black holes with non-spherically symmetric horizons or charged black holes. Additionally, it is natural to explore the Penrose inequality and the reverse isoperimetric inequality within cutoff holographic thermodynamics. The current study in AdS backgrounds can be extended to other contexts, such as cosmological holography Strominger (2001); Alishahiha et al. (2004); Gorbenko et al. (2019); Lewkowycz et al. (2020); Flauger et al. (2022); Silverstein and Torroba (2025); Aguilar-Gutierrez et al. (2025), Cauchy holography Araujo-Regado et al. (2023); Araujo-Regado (2022); Khan (2023); Soni and Wall (2024), and black holes with Lifshitz and hyperscaling violating asymptotics Cong et al. (2024). Furthermore, extensions to holographic thermodynamics with a negative deformation parameter $\lambda$ Guica and Monten (2021) can be considered. Investigating subleading orders in $1/N$ beyond the classical level presents an additional challenge for future work. Finally, the well-posedness of the Dirichlet boundary condition Liu et al. (2024, 2025) in finite-cutoff holography, the case of the mixed boundary conditions Guica and Monten (2021); He et al. (2022); He and Sun (2023), and other interpretation of holographic thermodynamics Mancilla (2024) warrant further study.
Acknowledgements—We are grateful to Rob Myers, David Kubizňák, Robie Hennigar, Xiaoyi Liu, Sergio E. Aguilar-Gutierrez, Song He, and Yuan Sun for helpful discussions. This work was supported by the Natural Sciences and Engineering Research Council of Canada and the National Natural Science Foundation of China (Grant No. 12365010). M. Z. and M. L. were supported by the China Scholarship Council Scholarship. W. T. was supported by a PhD grant from the China Scholarship Council. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.

### v2

@@ Abstract
The thermodynamics of AdS black holes has been extensively studied following the discovery of the Hawking-Page phase transition in Schwarzschild-AdS black holes Hawking and Page (1983b). Phase structures analogous to the van der Waals liquid–gas system have been identified for charged Reissner-Nordström AdS black holes, exhibiting first-order phase transitions and critical phenomena Chamblin et al. (1999a, b); Kubiznak and Mann (2012). Recently, significant advances have been achieved in AdS black hole thermodynamics by treating the cosmological constant as a variable, resulting in phenomena such as fluid/superfluid phase transitions Hennigar et al. (2017), black hole microstructures Wei et al. (2019), and more Kastor et al. (2009); Dolan (2011a, b); Lu et al. (2012); Xiao et al. (2024); Frassino et al. (2023). The study of thermodynamics within this extended phase space has developed into a subdiscipline known as black hole chemistry Kubiznak et al. (2017).
@@
on the conformal boundary at spatial infinity Ahmed et al. (2023a); Gong et al. (2023); Ahmed et al. (2023b). Here, $M,T,S,V,P$ respectively represent the bulk mass, temperature, entropy, thermodynamic volume, and thermodynamic pressure, while $\bar{E},\bar{T},\bar{S},\bar{V},\bar{P}$ denote their boundary counterparts. Note that there is an extra pair $(c,\mu)$ of conjugate variables, with $c$ being the central charge and $\mu$ its conjugate chemical potential, relative to the bulk. As pointed out in Visser (2022), the internal energy of the large-$N$ $SU(N)$ gauge theories with conformal symmetry depends not only on the extensive quantities $\bar{S}$ and $\bar{V}$, but also on the intensive quantity $c\sim N^{2}$. This is consistent with the scenario of working in the extended phase space, where $P=3/(8\pi\mathrm{G}\ell^{2})$ relates the AdS radius to the thermodynamic pressure, with Newton’s constant $\mathrm{G}$. Variation of the pressure is equivalent to varying central charge and volume of the CFT. The standard extrapolated holographic dictionary reads Witten (1998b); Gubser et al. (1996); Verlinde (2000); Savonije and Verlinde (2001)
@@
Finite-cutoff holographic thermodynamics: bulk side—The $T^{2}$ deformation of the three-dimensional CFT with large central charge is holographically dual to the SAdS black hole with a finite Dirichlet radial cutoff. In the framework of finite-cutoff holography, there is an agreement between the dressed energy spectrum (14) of the deformed field theory and the quasilocal energy of the SAdS black hole evaluated within the region $r<r_{c}$, given by Brown et al. (1994)
@@
Phase transitions—In the cutoff SAdS bulk, the free energy is defined as $G=\mathcal{E}-TS$; setting $M=0$ in (20) yields that of the cutoff AdS vacuum. As in the standard SAdS scenario, a Hawking–Page phase transition occurs: as the temperature increases, the system transits from the cutoff vacuum AdS to the cutoff SAdS black hole. On the deformed CFT side, comparing the free energy given by $\bar{G}=\bar{\mathcal{E}}-\bar{T}\bar{S}$ with that of the deformed vacuum reveals a confinement/deconfinement phase transition. Figure 1 shows the coexistence curve traced out by the phase transition points. Due to holographic duality, the $\bar{T}$ vs. $\bar{S}$ and $T/\omega$ vs. $S$ curves coincide and are plotted together. Varying $A_{c}$ or $\lambda$ amounts to changing the UV cutoff scale of the deformed CFT and selecting a family of $T^{2}$-deformed CFTs. This redefines the corresponding holographic RG flow starting from $A_{c}$ and also redefines the RG flow of the deformed CFT starting from a new UV scale. Note the deformation flow here is not the traditional RG flow; rather it defines a family of effective field theories without coarse-graining. The point $X_{0}$ marks the confinement-deconfinement transition point in the undeformed CFT or the Hawking–Page phase transition point in SAdS without a cutoff. As the deformation parameter $\lambda$ increases (reducing the cut-off $A_{c}$), the curve flows from $X_{0}\to X_{1}\to X_{2}\to X_{3}$, forming a teardrop-shaped Rupert coexistence curve. The region above the segment $X_{0}X_{1}X_{2}$ corresponds to the deconfinement (bulk black hole) phase, while the region below corresponds to the confinement (AdS vacuum) phase. As $\lambda$ increases further, the curve extends as entropy decreases back to $X_{0}$. The segment $X_{2}X_{0}$ separates the deconfinement (black hole) phase above from the confinement (AdS vacuum) phase below. The loop $X_{0}X_{1}X_{2}X_{0}$ encloses an island where the phase varies under the deformation flow, surrounded by an exterior region of fixed phase. Within the Rupert curve, up to three deformed CFTs or cut-off SAdS configurations may share the same phase transition temperature, with one deformed CFT (or cut-off SAdS) matching the seed CFT’s (or original SAdS’s) phase transition temperature and entropy. Note that according to (14), a shock singularity Smirnov and Zamolodchikov (2017) emerges for $\lambda>\pi\sqrt{c}\bar{V}^{3/2}/3\bar{S}^{3/2}$.
Figure 1: Coexistence curve of the boundary confinement/deconfinement and bulk Hawking-Page phase transitions with $c=4,\bar{V}=20$.
@@
Based on our results, several directions for future research can be pursued. Our analysis can be readily extended to holographic thermodynamics for a CFT in general dimensions, including $d=2$, where $T\bar{T}$ holographic thermodynamics can be constructed. The framework can also be applied to black holes with non-spherically symmetric horizons or charged black holes. Additionally, it is natural to explore the Penrose inequality Penrose (1973) and the reverse isoperimetric inequality Cvetic et al. (2011) within cutoff holographic thermodynamics. The current study in AdS backgrounds can be extended to other contexts, such as cosmological holography Strominger (2001); Alishahiha et al. (2004); Gorbenko et al. (2019); Lewkowycz et al. (2020); Flauger et al. (2022); Silverstein and Torroba (2025); Aguilar-Gutierrez et al. (2025), Cauchy holography Araujo-Regado et al. (2023); Araujo-Regado (2022); Khan (2023); Soni and Wall (2024), and black holes with Lifshitz and hyperscaling violating asymptotics Cong et al. (2024). Furthermore, extensions to holographic thermodynamics with a negative deformation parameter $\lambda$ Guica and Monten (2021) can be considered; the Hagedorn singularity that can be calculated via the torus partition function Gu et al. (2025) may arise and be related to the phase transition in the bulk. Investigating subleading orders in $1/N$ beyond the classical level presents an additional challenge for future work. Finally, the well-posedness of the Dirichlet boundary condition Liu et al. (2024, 2025) in finite-cutoff holography, the case of the mixed Guica and Monten (2021); He et al. (2022); He and Sun (2023) and more general Parvizi et al. (2025a, b) boundary conditions, and other relevant interpretations of black hole thermodynamics from holographic perfect fluid Mancilla (2024), particularly for the finite-cutoff scenario Tian et al. (2014), warrant further study.
Acknowledgements—We are grateful to Rob Myers, David Kubizňák, Robie Hennigar, Xiaoyi Liu, Sergio E. Aguilar-Gutierrez, Song He, Yuan Sun, and Yunfeng Jiang for helpful discussions. This work was supported by the Natural Sciences and Engineering Research Council of Canada and the National Natural Science Foundation of China (Grant No. 12365010). M. Z., W. T., and M. L. were supported by the China Scholarship Council Scholarship. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.